the percentage of oxygen in(

N

H

4

)

2

S

O

4

(NH4)2SO4






1
Expert's answer
2022-03-12T01:58:52-0500

Mr((NH4)2SO4) = (14 + 1×4)×2 + 32 + 16×4 = 132 g/mol

w(O) = (16×4 / 132) × 100% = 48.5%